OpenStudy (anonymous):

what is the value of x? 2xcubed-4=50

geerky42 (geerky42):

factor 2 out then divide both sides by 2. \(x^3-2 = 25\)

geerky42 (geerky42):

\(2x^3-4 = 2(x^3-2)\)

geerky42 (geerky42):

Then you add both sides by 2, then you have \(x^3 = 25+2 = 27\) Solving \(x^3=27\) should be easy